OpenStudy (anonymous):

When 1.470 × 10^3 J of heat energy is added to 36.9 g of hexane, C6H14, the temperature increases by 17.6 °C. Calculate the molar heat capacity of C6H14.

OpenStudy (gebooors):

Calculated molar mass of hexane, then amount of substance n. Molar heat capacity is n/ (energy * change of temperature), Unit is mol / J * degrees of C

OpenStudy (aaronq):

hm the formula above is incorrect, moles should be at the bottom because the units are \(energy/K*mol\). you wanna use: \(q=n*C*\Delta T\) and isolate C, specific heat capacity. you can substitute \(n=\dfrac{m}{M}\) into it, so you're left with: \(q=\dfrac{m}{M}*C*\Delta T \rightarrow C=\dfrac{q*M}{m*\Delta T}\)